A uniform wire of resistance 100 Ω is melted and recast in a wire of length double that of the original. What would be the resistance of the wire?

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. ℓ ‘ = 2 ℓ volume of the wire remains constant. A ℓ = A’ ℓ ‘ ⇒ A ℓ = A’ * 2 ℓ ⇒ A’ = A/2 f = Specific resistance R = fℓ/A : R’ = fℓ’/A’ 100 Ω = f2ℓ/A / 2 = 4fℓ/A = 4R ⇒ 4 * 100 Ω = 400 Ω